OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? −x + 2y = 9 4x + 5y = 9

OpenStudy (anonymous):

−x + 2y = 9 =>x=2y-9 4x + 5y = 9 => 4(2y-9)+5y=9 =>8y-36+5y=9 =>13y=45 y=45/13 find x by substituting y value in x

OpenStudy (anonymous):

in x equation

OpenStudy (anonymous):

whats the anser ?